# VER-Bench: Evaluating MLLMs on Reasoning with Fine-Grained Visual Evidence

### v1

@@ Abstract.
With the rapid development of MLLMs, evaluating their visual capabilities has become increasingly crucial. Current benchmarks primarily fall into two main types: basic perception benchmarks, which focus on local details but lack deep reasoning (e.g., ”what is in the image?”), and mainstream reasoning benchmarks, which concentrate on prominent image elements but may fail to assess subtle clues requiring intricate analysis. However, profound visual understanding and complex reasoning depend more on interpreting subtle, inconspicuous local details than on perceiving salient, macro-level objects. These details, though occupying minimal image area, often contain richer, more critical information for robust analysis. To bridge this gap, we introduce the VER-Bench, a novel framework to evaluate MLLMs’ ability to: 1) identify fine-grained visual clues, often occupying on average just 0.25% of the image area; 2) integrate these clues with world knowledge for complex reasoning. Comprising 374 carefully designed questions across Geospatial, Temporal, Situational, Intent, System State, and Symbolic reasoning, each question in VER-Bench is accompanied by structured evidence: visual clues and question-related reasoning derived from them. VER-Bench reveals current models’ limitations in extracting subtle visual evidence and constructing evidence-based arguments, highlighting the need to enhance models’ capabilities in fine-grained visual evidence extraction, integration, and reasoning for genuine visual understanding and human-like analysis. Dataset and additional materials are available https://github.com/verbta/ACMMM-25-Materials.
@@ 1. Introduction
This motivates our focus on small image regions, which often provide more decisive information than larger, salient areas. For example, in the right sub-figure of Fig.1, the distant statue in the bay serves as a distinctive landmark for precise localization, whereas the prominent skyline and building clusters offer limited diagnostic value. Similarly, while a cruise ship’s size or color provides ambiguous cues, fine details such as a hull number or unique logo enable definitive identification
@@ 2.2. MLLM Benchmarks
The rapid advancement of MLLMs has spurred the creation of comprehensive evaluation benchmarks (Duan et al., 2024; Fu et al., 2024). These benchmarks cover a wide range of capabilities, including real-world applications and visual perception (e.g., MME (Fu et al., 2024), MMBench (Liu et al., 2023a), MMVet (Yu et al., 2023)), integrated reasoning (e.g., MMStar (Chen et al., 2024a), MMMU (Yue et al., 2024)), and video understanding (e.g., MMBench-Video (Fang et al., 2025)). Specialized benchmarks also exist for tasks like chart interpretation (ChartQA (Masry et al., 2022)), document analysis (DocVQA (Mathew et al., 2021)), and mathematical reasoning (Yan et al., 2024). Evaluation methodologies primarily include Visual Question Answering (VQA), which allows open-ended responses but faces scoring objectivity issues (Papineni et al., 2002; Vedantam et al., 2015; Anderson et al., 2016), and Multiple Choice Questions (MCQ) (Schwenk et al., 2022), which are prevalent in leaderboards due to more reliable outcomes. However, existing MCQ benchmarks face several criticisms. Older benchmarks like MME (Fu et al., 2024) and POPE (Li et al., 2023) are often saturated, while newer, more challenging ones (e.g., MME-RealWorld (Zhang et al., 2024), RealWorldQA (XAI, 2024)) can suffer from ambiguity. Other concerns include answer leakage, partly addressed since MMStar (Chen et al., 2024a), and data redundancy or overlap (Zhang et al., 2025). Additionally, there is a recognized deficiency in benchmarks for specific industrial applications and non-English languages, with limited options like MME-Realworld-CN (Zhang et al., 2024) and MMBench-CN (Liu et al., 2023a).
@@ 3.1. Annotation Pipeline and Curation Process
The dataset is built via a multi-stage annotation pipeline with four key steps (as shown in Fig. 2) to generate high-quality visual evidence reasoning chains for real-world scenarios.
@@
Step 3: Answer and Visual Evidence Verification. This step ensured the dataset’s internal consistency and factual accuracy. We cross-verified answers, reasoning, and supporting visual evidence using multiple MLLMs (e.g., Qwen2.5-VL-72B and Claude-3.5-Sonnet). Only instances with consensus on the final answer and reasoning chain across all models were retained, minimizing model-specific biases and enhancing annotation reliability.
Step 4: Human Expert Review. Human experts conducted two review rounds to ensure accuracy and remove errors: 1. Visual Clue Verification: Experts confirmed the presence of each annotated visual clue in the image, eliminating any incorrect or hallucinated clues; 2. Reasoning Validation: Each reasoning step was reviewed for logical validity and alignment with real-world knowledge.
@@ 3.2. Dataset Overview
The dataset comprises 342 images and 374 Q&A pairs, with 614 visual clues annotated by 977 bounding boxes. These clues cover an average of 0.25% of the image area, making them ideal for fine-grained reasoning evaluation. VER-Bench is categorized into six tasks (Fig. 3), detailed below:
Geospatial Reasoning. Identifies geographic locations using visual clues like landmarks, scripts, symbols, and vehicle details. It requires global geography knowledge and cross-referencing multiple clues for accurate localization.
Temporal Reasoning. Infers time information (e.g., dates) from clues like clocks, text, seasonal signs, and event-specific elements.
Situational Adaptation Reasoning. Interprets complex scenes to infer situational contexts (e.g., events, activities) by understanding human behavior, spatial use, and environmental cues.
Intent Inference. Uncovers the underlying purpose of objects, designs, or behaviors, requiring semantic abstraction and contextual interpretation.
System State Understanding. Assesses the operational status of facilities or objects by analyzing visual signs like warning labels, crowd flow, or device status.
Symbolic Analysis. Recognizes and interprets cultural or historical symbols (e.g., emblems, icons) using both visual perception and deep cultural knowledge.
@@ 3.3. Evaluation Metric of VER-Bench
Visual reasoning differs from problems requiring strict sequential logic (e.g., math, programming). In visual reasoning, clue-reasoning pairs often lack strong sequential dependencies, and the order of identification typically does not affect the outcome. These pairs act as complementary ”evidence,” with more identified pairs enhancing answer reliability. Additionally, imperfect alignment between model-output and GT clues results in non-binary matching, complicating recall calculations.
@@ 4.1. Experiment Setups
We evaluated 12 open-source models, including Qwen2.5-VL (8B, 32B, 72B) (Bai et al., 2023), Qwen2-VL (8B) (Wang et al., 2024a), InternVL3 (8B, 38B, 78B) (Zhu et al., 2025), InternVL2.5 (8B) (Chen et al., 2024b), InternVL2 (8B) (Dong et al., 2024), MiniCPM-V-2.6 and MiniCPM-o-2.6 (Yao et al., 2024), and LLaVA-OneVision-72B (Li et al., 2024c). Additionally, we evaluated three closed-source models: Gemini-2.5-Pro-Preview (Team et al., 2023), GPT-4o (Achiam et al., 2023), and Claude-3.7-Sonnet (Anthropic, [n. d.]). Open-source models are tested on NVIDIA H20 GPUs using the MS-Swift framework, while closed-source models are evaluated via API calls. GPT-4o was used to score predictions against ground truth.
@@ 4.2. Quantitative Results
Closed-source models generally outperform open-source ones, with Gemini-2.5-Pro-Preview achieving the highest Average Accuracy of 76.8, surpassing GPT-4o by 12.5 points. Among open-source models, Qwen2.5-VL-32B achieved the top score of 60.8, outperforming larger models like Qwen2.5-VL-72B and InternVL3-78B. This is attributed to its reinforcement learning strategy, which enhances fine-grained visual analysis and instruction following. Qwen2.5-VL-72B closely follows at 59.6, while InternVL3-78B scored 56.9. LLaVA-OneVision-72B, despite its larger parameters, performs comparably to smaller Qwen-VL and InternVL models, likely due to its focus on cross-modal migration to multi-images and videos rather than single-image understanding. The edge models MiniCPM-V 2.6 and MiniCPM-o 2.6 show strong performance for their size, surpassing InternVL2-8B and InternVL2.5-8B.
@@ 4.2.3. Reasoning Quality
measures the semantic consistency between the reasoning generated by the model and the GT reasoning. A high value signifies that the model’s logical deduction process closely aligns with the expected correct reasoning process. Gemini-2.5-Pro-Preview leads with 69.4 points, demonstrating its ability not only to identify clues correctly but also to explain how these clues lead to conclusions in a manner akin to human standards. GPT-4o (54.5) and Claude-3.7-Sonnet (43.4) also exhibit strong reasoning quality. Among open-source models, Qwen2.5-VL-72B (46.6), InternVL3-78B (45.7), and Qwen2.5-VL-32B (45.6) stand out, indicating they excel at simulating correct reasoning paths compared to other open-source counterparts. In contrast, models such as InternVL2.5-8B (25.5) and InternVL2-8B (22.2) score lower. This suggests their reasoning processes may contain logical flaws, insufficient explanations, or significant semantic deviations from GT reasoning—even if they identify some correct visual clues.
@@ 4.2.4. Evidence-to-Answer Relevance
Across all metrics, the average score for is significantly higher than others, with even the lowest score exceeding 76. This indicates that even if a model’s final answer () or reasoning process () deviates from the GT, it can still construct internally self-consistent ”evidence-reasoning-answer”. For instance, Qwen2.5-VL-72B outperformed closed-source models like Gemini-2.5-Pro-Preview in , highlighting its potential for generating coherent explanations. Conversely, lower scores, such as MiniCPM-o 2.6 (76.0) and InternVL2.5-8B (77.5), suggest that these models’ reasoning steps may be more divergent or contain redundant information with limited support for their conclusions. is crucial for evaluating the coherence and persuasiveness of model outputs. A high score signifies a more rigorous and focused argumentation process.

### v2

@@ Abstract.
With the rapid development of MLLMs, evaluating their visual capabilities has become increasingly crucial. Current benchmarks primarily fall into two main types: basic perception benchmarks, which focus on local details but lack deep reasoning (e.g., ”what is in the image?”), and mainstream reasoning benchmarks, which concentrate on prominent image elements but may fail to assess subtle clues requiring intricate analysis. However, profound visual understanding and complex reasoning depend more on interpreting subtle, inconspicuous local details than on perceiving salient, macro-level objects. These details, though occupying minimal image area, often contain richer, more critical information for robust analysis. To bridge this gap, we introduce the VER-Bench, a novel framework to evaluate MLLMs’ ability to: 1) identify fine-grained visual clues, often occupying, on average, just 0.25% of the image area; 2) integrate these clues with world knowledge for complex reasoning. Comprising 374 carefully designed questions across Geospatial, Temporal, Situational, Intent, System State, and Symbolic reasoning, each question in VER-Bench is accompanied by structured evidence: visual clues and question-related reasoning derived from them. VER-Bench reveals current models’ limitations in extracting subtle visual evidence and constructing evidence-based reasoning chains, highlighting the need to enhance models’ capabilities in fine-grained visual evidence extraction, integration, and reasoning for genuine visual understanding and human-like analysis. The dataset is available at https://github.com/verbta/ACMMM-25-Materials.
@@ 1. Introduction
This motivates our focus on small image regions, which often provide more decisive information than larger, salient areas. For example, in the right sub-figure of Fig.1, the distant statue in the bay serves as a distinctive landmark for precise localization, whereas the prominent skyline and building clusters offer limited diagnostic value. Similarly, while a cruise ship’s size or color provides ambiguous cues, fine details such as a hull number or unique logo enable definitive identification.
@@ 2.2. MLLM Benchmarks
The rapid advancement of MLLMs has spurred the creation of comprehensive evaluation benchmarks (Duan et al., 2024; Fu et al., 2024). These benchmarks cover a wide range of capabilities, including real-world applications and visual perception (e.g., MME (Fu et al., 2024), MMBench (Liu et al., 2023a), MMVet (Yu et al., 2023)), integrated reasoning (e.g., MMStar (Chen et al., 2024a), MMMU (Yue et al., 2024)), and video understanding (e.g., MMBench-Video (Fang et al., 2025)). Specialized benchmarks also exist for tasks like chart interpretation (ChartQA (Masry et al., 2022)), document analysis (DocVQA (Mathew et al., 2021)), and mathematical reasoning (Yan et al., 2024). Evaluation methods primarily include Visual Question Answering (VQA), which allows open-ended responses but faces scoring objectivity issues (Papineni et al., 2002; Vedantam et al., 2015; Anderson et al., 2016), and Multiple Choice Questions (MCQ) (Schwenk et al., 2022), which are prevalent in leaderboards due to more reliable outcomes. However, existing MCQ benchmarks face several criticisms. Older benchmarks like MME (Fu et al., 2024) and POPE (Li et al., 2023) are often saturated, while newer, more challenging ones (e.g., MME-RealWorld (Zhang et al., 2024), RealWorldQA (XAI, 2024)) can suffer from ambiguity. Other limitations include answer leakage (mitigated post-MMStar (Chen et al., 2024a)) and data redundancy (Zhang et al., 2025). Additionally, there is a recognized deficiency in benchmarks for specific industrial applications and non-English languages, with limited options like MME-RealWorld-CN (Zhang et al., 2024) and MMBench-CN (Liu et al., 2023a).
@@ 3.1. Annotation Pipeline and Curation Process
We construct the dataset through a four-stage annotation pipeline (as shown in Fig. 2) to generate high-quality visual evidence reasoning chains for real-world scenarios.
@@
Step 3: Answer and Visual Evidence Verification. This step ensures the dataset’s internal consistency and factual accuracy. We cross-verified answers, reasoning, and supporting visual evidence using multiple MLLMs (e.g., Qwen2.5-VL-72B, Claude-3.5-Sonnet). Only instances with consensus on the final answer and reasoning chain across all models were retained, to minimize model-specific biases and enhancing annotation reliability.
Step 4: Human Expert Review. In two review rounds, human experts ensure accuracy and remove errors: 1. Visual Clue Verification: Experts confirmed the presence of each annotated visual clue in the image, eliminating any incorrect or hallucinated clues; 2. Reasoning Validation: Each reasoning step was reviewed for logical validity and alignment with real-world knowledge.
@@ 3.2. Dataset Overview
The dataset comprises 342 images and 374 Q&A pairs, with 614 visual clues annotated with 977 bounding boxes. These clues cover an average of 0.25% of the image area, making them ideal for fine-grained reasoning evaluation. VER-Bench is categorized into six tasks (Fig. 3), detailed below:
Geospatial Reasoning. This task identifies geographic locations using visual clues like landmarks, scripts, symbols, and vehicle details. It requires global geography knowledge and cross-referencing multiple clues for accurate localization.
Temporal Reasoning. This task infers time information (e.g., dates) from clues like clocks, text, seasonal signs, and event-specific elements.
Situational Adaptation Reasoning. This task interprets complex scenes to infer situational contexts (e.g., events, activities) by understanding human behavior, spatial use, and environmental cues.
Intent Inference. This task uncovers the underlying purpose of objects, designs, or behaviors, requiring semantic abstraction and contextual interpretation.
System State Understanding. This task assesses the operational status of facilities or objects by analyzing visual signs like warning labels, crowd flow, or device status.
Symbolic Analysis. This task recognizes and interprets cultural or historical symbols (e.g., emblems, icons) using both visual perception and deep cultural knowledge.
@@ 3.3. Evaluation Metric of VER-Bench
Visual reasoning differs from problems requiring strict sequential logic (e.g., math, programming). In visual reasoning, clue-reasoning pairs often lack strong sequential dependencies, and the order of identification typically does not affect the outcome. These evidence pairs are complementary: identifying more pairs improves answer reliability. Additionally, imperfect alignment between model-output and GT clues results in non-binary matching, complicating recall calculations.
@@ 4.1. Experiment Setups
We evaluate 12 open-source models, including Qwen2.5-VL (8B, 32B, 72B) (Bai et al., 2023), Qwen2-VL (8B) (Wang et al., 2024a), InternVL3 (8B, 38B, 78B) (Zhu et al., 2025), InternVL2.5 (8B) (Chen et al., 2024b), InternVL2 (8B) (Dong et al., 2024), MiniCPM-V-2.6, MiniCPM-o-2.6 (Yao et al., 2024), and LLaVA-OneVision-72B (Li et al., 2024c). Additionally, we evaluate three closed-source models: Gemini-2.5-Pro-Preview (Team et al., 2023), GPT-4o (Achiam et al., 2023), and Claude-3.7-Sonnet (Anthropic, [n. d.]). Open-source models are tested on NVIDIA H20 GPUs using the MS-Swift framework, while closed-source models are evaluated via API calls. GPT-4o is used to score predictions against ground truth.
@@ 4.2. Quantitative Results
Closed-source models lead in performance, with Gemini-2.5-Pro-Preview scoring 76.8 average accuracy, surpassing GPT-4o by 12.5 points. Among open-source models, Qwen2.5-VL-32B tops at 60.8, outperforming larger variants due to its reinforcement learning strategy that improves visual detail understanding and instruction following. Qwen2.5-VL-72B (59.6) and InternVL3-78B (56.9) trail closely. LLaVA-OneVision-72B performs on par with smaller models, likely because it prioritizes multi-image and video tasks over single-image reasoning. Lightweight models MiniCPM-V 2.6 and MiniCPM-o 2.6 also punch above their weight, outperforming larger 8B-class models.
@@ 4.2.3. Reasoning Quality
measures how semantically consistent the model’s reasoning is with the ground truth. A higher score indicates closer alignment with correct logical deduction. Gemini-2.5-Pro-Preview leads at 69.4, showing strong human-like reasoning. GPT-4o (54.5) and Claude-3.7-Sonnet (43.4) also perform well. Among open-source models, Qwen2.5-VL-72B (46.6), InternVL3-78B (45.7), and Qwen2.5-VL-32B (45.6) stand out, demonstrating effective reasoning. In contrast, models such as InternVL2.5-8B (25.5) and InternVL2-8B (22.2) score lower. This suggests their reasoning processes may contain logical flaws, insufficient explanations, or significant semantic deviations from GT reasoning—even if they identify some correct visual clues.
@@ 4.2.4. Evidence-to-Answer Relevance
Across all metrics, the average score for is significantly higher than others, with even the lowest score exceeding 76. This indicates that even if a model’s final answer () or reasoning process () deviates from the GT, it can still construct internally self-consistent ”evidence-reasoning-answer”. For instance, Qwen2.5-VL-72B outperformed closed-source models like Gemini-2.5-Pro-Preview in , highlighting its potential for generating coherent explanations. Conversely, lower scores, such as MiniCPM-o 2.6 (76.0) and InternVL2.5-8B (77.5), suggest that these models’ reasoning steps may be more divergent or contain redundant information with limited support for their conclusions. is crucial for evaluating the coherence and persuasiveness of model outputs. A high score signifies a more rigorous and focused reasoning process.